An apple sells for 25 cents and a peach sells for 15 cents . A total of 10 pieces of fruit were sold for a total cost of $ 2.10. How many apples were sold
Answer by lwsshak3(11628)

An apple sells for 25 cents and a peach sells for 15 cents . A total of 10 pieces of fruit were sold for a total cost of $ 2.10. How many apples were sold
**
let x=number of apples sold
10-x=nmber of peaches sold
..
.25x+.15(10-x)=2.10
.25x+1.5-.15x=2.10
.10x=.60
x=6
10-x=4
ans:
number of apples sold=6
nmber of peaches sold=4